OpenStudy (anonymous):

Here is a question in which I know what to do, but I do not get the right answer. I get very confused, so I am wondering if any of you can tell me what I am doing wrong. Here is the problem: the coordinates of a isosceles trapezoid ABCD are A(-r,s) B(r,s) C(t,0) and D(t,0) Right, so my plan is: 1. calculate distance between AC 2. Calculate distance between DB 3. Compare the distances But once I get to it, all the letters confuse me and what I get is really out of place, how would you guys do this?

OpenStudy (chaise):

A-C (-r, s)-(t,0)={(-r-t),s} B-D (r, s)-(t, 0)={(r-t), s)} Does this help at all? I am pretty sure from this we can conclude that the distances are equal, because the fact that the (r-t) might be negative just shows a reflection in sides, for example the length AC is reflected BD.. if that sort of makes sense? Try drawing this shape to get the idea if your still stuck!

OpenStudy (amistre64):

what are you trying to find with this iso trap?

OpenStudy (amistre64):

and is C equal to D?

OpenStudy (anonymous):

I am using the distance formula for this but whenever I do it, I get something that I don't know if it is right

OpenStudy (amistre64):

i dont know if the distance formula even needs to be used simply because i do not know what the question is asking for. why is the distance formula needed in this situation?

OpenStudy (amistre64):

are you trying to determine if the diagonals are equal?

OpenStudy (amistre64):

A(-r,s) to C(t,0) B(r,s) to D(t,0) ^^ is one of these a -t?

OpenStudy (anonymous):

You need the distance formula to compare them, cause it asks you to prove it

OpenStudy (phi):

First, you have to fix this C(t,0) and D(t,0) the two points should not have the same coordinates

OpenStudy (anonymous):

Hmmm then I have to ask the teacher again...

OpenStudy (phi):

I think amistre has it right, and you have a typo. one should be (-t,0)

OpenStudy (phi):

if we go on the assumption that C is at (-t,0) then you can use the distance formula remember it is \[ d= \sqrt{ (y_1 - y_0)^2 + (x_1 - x_0)^2 } \]

OpenStudy (anonymous):

hmm no, I will try it that way but the problem does not say C is that, maybe a study guide typo.

OpenStudy (anonymous):

Hmm guys, I don't know, im usually good at distance formulas but add the letters and I am sooo bad at it .